OpenStudy (asylum15):

- Can sin^3x be divided by sinx?

OpenStudy (asylum15):

\[\int\limits (u)^4 . \sin^3x . \frac{ du }{ sinx }\] Can I divid the sin^3x by sinx under the line to cancel?

OpenStudy (anonymous):

Yes, just a simplify step!

OpenStudy (asylum15):

Becomes sin^2x ?

OpenStudy (anonymous):

Yes, sin²x

OpenStudy (asylum15):

<3!

OpenStudy (anonymous):

What's the original problem?